5 workers can make 10 toys in 6 days ,6 hours per day.12 workers can make 16 toys in how many days ,8 hours per day?

5 workers can  make 10 toys in 6 days ,6 hours per day.12 workers can make 16 toys in how many days ,8 hours per day?

Grade:12th pass

2 Answers

hari krishn
11 years ago

Ans: 3 days

If all the workers have the same potential in their work....

1 worker can make 2 toys in 6*6=36 hours....

12 workers can make 24 toys in 36 hours

so, 12 workers can make 16 toys in 16*36/24 = 24 hours....

as such 8 hours of working per day, 12 workers can make 16 toys in 3 days.............

KUMAR PRASSU
11 years ago

From given question total houres required to make 10 toys =5x6x6 = 180 hrs

                                                   and for making 16 toys =180x16/10= 18x16hrs

let no.of days be ''d''

ie.,18x16 hrs are required therefore number of houres =>18x16=12xdx8=>d=3

there fore 3 days are requiredLaughing